Grade upto college level Physical Chemistry

An athlete is given 100 gm of glucose (C6H12O6) of energy equivalent to 1560 kJ. He utilizes 50 percent fo this gained energy in the event. In order to avoids storage of energy in the body, calculate the weight of water he would need to perspire. The enthalpy of evaporation of water is 44 kJ/mole.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
100 g of glucose = 1560 kJ
Energy utilized in body = 50/100 * 1560 = 780 kJ
Energy left unutilized in body = 1560 – 780 = 780 kJ
Energy to be given out = 1560 – 780 = 780 kJ
Enthalpy of evaporation of water = 44 kJ/mole = 44 kJ/18 g
Of water [1 mole H2O = 18g water]
Hence amount water to be perspired to avoid storage of energy = 18/44 * 780 = 319.1 g

Thanks
jitender
askIITians Faculty